# Adventures in Algebraic Geometry

### v1

@@
[…]
@@
 2.12.12.1 There are a lot of horizontal lines in these notes to remind the reader to take a sip of their drink of choice.
@@
has no zeros. Here $(I)$ denotes the ideal in the ring ${\mathbb{C}}[x_{1},\ldots,x_{n},t]$ generated by $I$. Indeed, any point $(a_{1},\ldots,a_{n},b)$ in the set of zeros of $\hat{I}$ must have $(a_{1},\ldots,a_{n})$ in the set of zeros of $I$. Then $ft-1$ takes value $0\,b-1\neq 0$ on $(a_{1},\ldots,a_{n},b)$, contradiction.
By Corollary 2.6, the ideal $\hat{I}$ is the whole ring, and in particular it contains $1$.2.22.22.2 It is a common trick to replace the statement that an ideal is equal to a ring by the statement that this ideal contains $1$. If you study commutative algebra long enough, you will get sick of it. In other words, we have in ${\mathbb{C}}[x_{1},\ldots,x_{n},t]$
@@
Every closed subset $Z(I)$ can be uniquely written as a disjoint union of irreducible closed subsets of $Z(I)$ none of which is contained in another one.
@@
In the existence direction, we start with $Z(I)$ and keep decomposing it into disjoint unions of smaller subsets. If we could keep doing this indefinitely, we would get a decreasing chain of closed subsets which are properly contained in one another. This would correspond to an increasing chain of ideals of ${\mathbb{C}}[x_{1},\ldots,x_{n}]$, in contradiction with it being Noetherian. Thus, we can write $Z(I)$ as a finite union of irreducible subsets, and it remains to remove the subsets that are contained in others.
@@
[…]
@@
It remains to argue that $K$ and $C$ are finitely generated modules over ${\mathbb{C}}[x_{0},\ldots,x_{n-1}]$ and apply induction on $n$ (with $n=0$ case just graded finite dimensional vector spaces) to see that $f(t)$ is a rational function with denominator that is a power of $(1-t)$. The relation to $\dim Z(I)$ is significantly more difficult to prove and the interested reader is referred to [2] or [12].
@@
To find a submodule of $M$ isomorphic to $(A/p)(d)$, let us ask ourselves what this means. Such isomorphism would send $1\in(A/p)(d)$ to some element $m\in M$. The degree of $1$ in $(A/p)(d)$ is, by the definition of the grading shift,5.15.15.1 Never trust any signs in any mathematical text ever! equal to $(-d)$, and the ideal $p$ can be recovered as
$p={\rm Ann}(m):=\{a\in A,{\rm~{}such~{}that~{}}a\hskip 1.0ptm=0\}.$
@@
So now we are simply looking for a nonzero homogeneous element $m\in M$ whose annihilator is prime. As with many things in mathematics, we will try to find such $m$ by imposing on it some maximum or minimum property. As it happens, the right thing to consider is nonzero homogeneous $m$ such that ${\rm Ann}(m)$ is maximum among all such annihilators, in the sense that there are no nonzero homogeneous $m^{\prime}$ with ${\rm Ann}(m^{\prime})\supsetneq{\rm Ann}(m)$. Existence of such $m$ is assured by $A$ being Noetherian. Now suppose that the ideal ${\rm Ann}(m)$ is not prime. It means that there exist $a,b\in A$ with $a\hskip 1.0ptb\hskip 1.0ptm=0$, $a\hskip 1.0ptm\neq 0$, $b\hskip 1.0ptm\neq 0$. Then ${\rm Ann}(b\hskip 1.0ptm)$ contains ${\rm Ann}(m)$ and it is strictly larger because it contains $a$ while ${\rm Ann}(m)$ does not. This finishes the proof. ∎
@@
$0\to{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1})(-\deg f_{2})\stackrel{{\scriptstyle f_{2}}}{{\longrightarrow}}{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1})\to{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1},f_{2})\to 0$
@@
In particular, $S$ contains $x_{0}$, and if we localize ${\mathbb{C}}[x_{0},x_{1},x_{2}]$ at it first we obtain ${\mathbb{C}}[\frac{x_{1}}{x_{0}},\frac{x_{2}}{x_{0}},x_{0}^{\pm 1}]$ (Laurent polynomials in $x_{0}$ over the polynomial ring ${\mathbb{C}}[y_{1},y_{2}]$ for $y_{i}=\frac{x_{i}}{x_{0}}$). When we localize further by the image of $S$ in ${\mathbb{C}}[\frac{x_{1}}{x_{0}},\frac{x_{2}}{x_{0}},x_{0}^{\pm 1}]$ we end up with the ${\mathbb{Z}}$-graded ring $R[x_{0}^{\pm 1}]$ where the ring $R$ is defined in (4.2). If we localize the module $M$, we end up with $(R/(\tilde{f}_{1},\tilde{f}_{2}))[x_{0}^{\pm 1}]$ where $\tilde{f}_{i}=\frac{f_{i}}{x_{0}^{\deg f_{i}}}$ is the dehomogenization of $f_{i}$. We can then take degree 00 part of this space to get exactly $R/(\tilde{f}_{1},\tilde{f}_{2})$ that was used to define the multiplicity of the intersection at the point $(1:0:0)$.
So we will take the filtration of $M$ from Proposition 5.1 and observe that the operations of taking localization in $S$ and degree 00 parts preserves the filtration property, with the associated graded objects changing accordingly. If we localize $M={\mathbb{C}}[x_{0},x_{1},x_{2}]/(x_{0},x_{1},x_{2})$ in $S$ we get 00, because $x_{0}$ both annihilates $M_{S}$ and is invertible. Similarly, localization of ${\mathbb{C}}[x_{0},x_{1},x_{2}]/p$ for a prime ideal that corresponds to a point $p$ other than $(1:0:0)$ is zero because there is an element in $S$ which is zero at $p$ but not at $(1:0:0)$ and is therefore both zero and invertible on the localization. Finally, $({\mathbb{C}}[x_{0},x_{1},x_{2}]/(x_{1},x_{2}))_{S}={\mathbb{C}}[x_{0}^{\pm 1}]$ and the degree zero of any grading shift of it is a dimension one vector space over ${\mathbb{C}}$. Thus, the number of occurrences of the ideal of $(1:0:0)$ in the filtration is exactly equal to the dimension of $R/(\tilde{f}_{1},\tilde{f}_{2})$ which is the multiplicity of the intersection at this point by (4.3). The same is true for all other points in ${\mathbb{C}}{\mathbb{P}}^{2}$, which finishes the proof of Bezout’s theorem.
@@
$\begin{array}[]{c}\pi^{-1}(U)\hskip 5.0pt\cong\hskip 5.0pt{\mathbb{C}}^{r}\times U\\ \searrow\hskip 20.0pt\swarrow\\ U\end{array}$ (6.4)
@@
Picard group ${\rm Pic}(X)$ is the (abelian) group of line bundles on $X$, up to isomorphism (as line bundles). The product is $\otimes$, the identity is the trivial line bundle ×${\mathbb{C}}\times X\to X$, and the inverse is $W\mapsto W^{\vee}$.
@@
The simplest example of a section is the zero section. This means that in every fiber we pick the well-defined element 00 of the corresponding vector space structure. For some line bundles this is the only (holomorphic) section.
@@
  We will also introduce some natural bundles on smooth manifolds which will be used later. Namely, for any smooth manifold $X$ we can construct its holomorphic tangent bundle $TX\to X$. Sections of this bundle on $U$ are the same as vector fields on $U$. Its dual $TX^{\vee}\to X$ is called the cotangent bundle. Sections of the cotangent bundle over $U$ are precisely the holomorphic $1$-forms on $U$6.36.36.3 When I was learning calculus, the symbol $dx$ really confused me. Was it a function of two variables? Of one variable? I only found peace when it became clear that $dx$ is a section of the cotangent bundle.
@@
Exercise 3. Verify that sections of the line bundle ×${\mathbb{C}}\times X\to X$ are in natural bijection to holomorphic functions on $X$.
@@
and note that sections $x_{i}$ of $\mathcal{O}(1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ pull back to holomorphic sections $s_{i}=f^{*}x_{i}$ of $f^{*}\mathcal{O}(1)\to X$. An important observation is that for every point $p\in X$ at least one of the sections $f^{*}x_{i}$ does not equal zero in the fiber of $f^{*}\mathcal{O}(1)\to X$7.17.17.1 We can not make sense of a value of a section, because identification of the fiber with ${\mathbb{C}}$ is not canonical, but we can ask whether the value is zero, since the fiber has a natural vector space structure. Indeed, the fiber of $f^{*}\mathcal{O}(1)\to X$ at $p$ is the fiber of $\mathcal{O}(1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ at $f(p)$ and we know that not all $x_{i}$ are zero at $f(p)$.
@@
Note that this map is well-defined and holomorphic. The easiest way to see both is to appeal to a local biholomorphism with ×${\mathbb{C}}\times U$. While not unique, it is defined up to a holomorphic invertible scaling of the fiber, so all $s_{i}(p)$ are multiplied by the same factor.
@@
As with the Weil divisors, there is a concept of a principal Cartier divisor. It is simply given by $(X,f)$ for a rational function $f$. Clearly, these are identified with principal Weil divisors. In short, we have the following diagram of abelian groups8.68.68.6 I have debated whether to write $1$ or 00 on the left, but chose $1$ to honor the multiplicative nature of ${\rm InvRat}(X)$.
$\begin{array}[]{c}1\to{\rm InvRat}(X)\to{\rm Rat}(X)^{*}\to{\rm CartierDiv}(X)\to{\rm CaCl}(X)\to 0\\ \Arrowvert\hskip 60.0pt\Arrowvert\hskip 70.0pt\downarrow\hskip 60.0pt\downarrow\\ 1\to{\rm InvRat}(X)\to{\rm Rat}(X)^{*}\to\hskip 7.0pt{\rm WeilDiv}(X)\hskip 6.0pt\to\hskip 7.0pt{\rm Cl}(X)\hskip 7.0pt\to 0\end{array}$ (8.2)
@@
Show that the ideal $I=(x_{1},x_{3})\subset R$ is not principal.8.88.88.8 This is an infinitesimal version of the statement that the Weil divisor on the surface $x_{1}x_{2}=x_{3}^{2}$ given by $[x_{1}=x_{3}=0]$ can not be locally given as ${\rm div}f$. Hint: Look at the dimension of $I/(x_{1},x_{2},x_{3})I$.
@@
Let $X$ be a smooth complex manifold. Recall from Section 6 that a line bundle $\pi:L\to X$ is a holomorphic map from a smooth complex manifold $L$, with the fibers $\pi^{-1}(x)$ given a structure of a one-dimensional complex vector space. The simplest example is that of a trivial line bundle ×${\mathbb{C}}\times X\to X$ and all line bundles are glued from ×${\mathbb{C}}\times U\to U$ for some open subsets $U\subseteq X$.
Recall that holomorphic sections of $\pi:L\to X$ are holomorphic maps $s:X\to L$ such that $\pi\circ s={\rm id}_{X}$. If $L={\mathbb{C}}\times X\to X$ is a trivial line bundle, then holomorphic sections of it are in bijection with holomorphic functions $X\to{\mathbb{C}}$. Similarly, we can talk about meromorphic sections of line bundles. Indeed, locally every line bundle is (non-canonically) isomorphic to ×${\mathbb{C}}\times U\to U$ and we would be looking at meromorphic functions on $U$ (ratios of two holomorphic functions). If we change to another identification of $\pi^{-1}U$ with ×${\mathbb{C}}\times U$, we simply multiply by an invertible holomorphic function, so meromorphicity of sections is well defined.
@@
The identity of the group structure is given by the trivial line bundle ×${\mathbb{C}}\times X\to X$ with the section $x\mapsto(1,x)$, and it is easy to check group properties. When checking that the map is a homomorphism, we can work on a cover $X=\bigcup_{\alpha}U_{\alpha}$ for which both $L_{1}$ and $L_{2}$ have trivializations. Then tensor product of sections gives rise to the product of the corresponding functions $f_{\alpha,1}$ and $f_{\alpha,2}$, which corresponds to the product of the Cartier divisors. ∎
@@
Suppose we are given a Cartier divisor $(U_{\alpha},f_{\alpha})_{\alpha\in I}$. Take a copy of ×${\mathbb{C}}\times U_{\alpha}$ for each $\alpha$ and glue them together by the following identification. We consider the equivalence relation of the set $\bigsqcup_{\alpha}({\mathbb{C}}\times U_{\alpha})$ with
(×$({\mathbb{C}}\times U_{\alpha})\ni(t,x)\sim\left(\frac{f_{\beta}(x)}{f_{\alpha}(x)}\,t,x\right)\in({\mathbb{C}}\times U_{\beta})$
@@
For a rational function $f$ on $X$, the principal Cartier divisor ${\rm div}f$ corresponds to the trivial line bundle ×${\mathbb{C}}\times X\to X$ with the rational section identified with $f$, i.e.
@@
We can easily verify that the above description fits the definition of Cartier divisor. For example, the ratio $\frac{f_{1}}{f_{2}}=\frac{x_{2}}{x_{1}}$ is an invertible holomorphic function on $U_{1}\cap U_{2}$. The corresponding Weil divisor is built from the zeros and poles of $f_{i}$ and is the line [$[\{x_{0}=0\}]$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, with coefficient $1$.
@@
This gives us $f_{1}=\left(\frac{x_{0}}{x_{1}}\right)^{-3}$ (signs, or more generally invertible functions do not matter) and similarly $f_{2}=\left(\frac{x_{0}}{x_{2}}\right)^{-3}$. Then the corresponding Weil divisor is $-3[\{x_{0}=0\}]$, and therefore the canonical line bundle over ${\mathbb{C}}{\mathbb{P}}^{2}$ is isomorphic to $\mathcal{O}(-3)$.
@@
$\left\{\begin{array}[]{ll}y^{2}-(x-\alpha_{1})(x-\alpha_{2})(x-\alpha_{3})=0,&\\ -(x-\alpha_{2})(x-\alpha_{3})-(x-\alpha_{1})(x-\alpha_{2})-(x-\alpha_{1})(x-\alpha_{2})=0,&\\ 2y=0.&\end{array}\right.$
@@
[…]
@@
for the newly defined addition operation $+$. We claim that it gives $E$ a group structure, which is of course abelian by construction, since lines $AB$ and $BA$ are the same.
@@
[…]
@@
[…]
@@
We know that the residue at $z=0$ must be 00, so our first attempt at finding such $f$ is
@@
Great, this looks periodic. But, we have a slight problem – the series does not have absolute convergence. Indeed, the number of lattice points $l$ of size approximately $k$ is on the order of $c\,k$ for some constant $c$, so we have roughly $c\,k$ terms of size $\frac{1}{k^{2}}$, and the harmonic series diverges (although rather slowly).
@@
In the last equality we used that the change of variables $l^{\prime}\to-m-l^{\prime}$ sends $S=\sum_{l^{\prime}\neq-m,0}\Big{(}\frac{1}{(l^{\prime})^{2}}-\frac{1}{(l^{\prime}+m)^{2}}\Big{)}$ to $(-S)$, which implies $S=0$.
@@
For every $a\not\in\frac{1}{2}L$ the even elliptic function $\mathcal{P}(z)-\mathcal{P}(a)$ has zeros at $a\,{\rm mod}\,L$ and $-a\,{\rm mod}\,L$ with multiplicity one, and no other zeros. For $a\in\frac{1}{2}L\setminus L$ the function $\mathcal{P}(z)-\mathcal{P}(a)$ has zeros of order two at $a\,{\rm mod}\,L$ and no other zeros.
@@
The function $\mathcal{P}(z)-\mathcal{P}(a)$ has a pole of order two at $L$ and no other poles. Therefore, by Propositions 11.3 and 11.4 it has two zeros (with multiplicity, up to $L$) which add up to $0\,{\rm mod}\,L$. One of these zeros is $a$, so the other is $(-a)$. The situation with $2a\in L$ is special, because then we have $a=-a\,{\rm mod}\,L$ and a double zero. ∎

### v2

@@
$\bullet$~~ $xy-1=0\mathrm{~in~}{\mathbb{C}}^{2}.\hskip 80.0pt~~\bullet~~x^{2}-y^{2}=1\mathrm{~in~}{\mathbb{Q}}^{2}.$
$\bullet$~~ $\left\{\begin{array}[]{r}xt-yz=0\\ xz-y^{2}=0\\ yt-z^{2}=0\end{array}\right.\mathrm{~in~}{\mathbb{C}}^{2}.\hskip 50.0pt~~\bullet~~y^{2}-y=x^{3}-x\mathrm{~in~}{\mathbb{Q}}^{2},{\mathbb{R}}^{2},\mathrm{~or~}{\mathbb{C}}^{2}.$
$\bullet$~~ $x^{2}=y^{3}\mathrm{~in~}{\mathbb{C}}^{2}.\hskip 94.0pt~~\bullet~~\left\{\begin{array}[]{l}\det A=0\\ \mathrm{tr}\,A=0\end{array}\right.\mathrm{~in~}{\mathbb{C}}^{n^{2}}.$
$\bullet$~~ $\mathrm{all~}3\times 3\mathrm{~minors~of~skew-symmetric~}n\times n\mathrm{~matrices~are~zero,}\mathrm{~in~}{\mathbb{C}}^{\frac{n(n-1)}{2}}.$
@@
 2.12.12.1There are a lot of horizontal lines in these notes to remind the reader to take a sip of their drink of choice.
@@
has no zeros. Here $(I)$ denotes the ideal in the ring ${\mathbb{C}}[x_{1},\ldots,x_{n},t]$ generated by $I$. Indeed, any point $(a_{1},\ldots,a_{n},b)$ in the set of zeros of $\hat{I}$ must have $(a_{1},\ldots,a_{n})$ in the set of zeros of $I$. Then ft1ft-1 takes value $0\,b-1\neq 0$ on $(a_{1},\ldots,a_{n},b)$, contradiction.
By Corollary 2.6, the ideal $\hat{I}$ is the whole ring, and in particular it contains 11.2.22.22.2It is a common trick to replace the statement that an ideal is equal to a ring by the statement that this ideal contains $1$. If you study commutative algebra long enough, you will get sick of it. In other words, we have in ${\mathbb{C}}[x_{1},\ldots,x_{n},t]$
@@
Every closed subset $Z(I)$ can be uniquely written as a union of irreducible closed subsets of $Z(I)$ none of which is contained in another one.
@@
In the existence direction, we start with $Z(I)$ and keep decomposing it into unions of smaller subsets. If we could keep doing this indefinitely, we would get a decreasing chain of closed subsets which are properly contained in one another. This would correspond to an increasing chain of ideals of ${\mathbb{C}}[x_{1},\ldots,x_{n}]$, in contradiction with it being Noetherian. Thus, we can write $Z(I)$ as a finite union of irreducible subsets, and it remains to remove the subsets that are contained in others.
@@
$\bullet$ $\{x_{0}x_{2}-x_{1}^{2}=0\}\subset{\mathbb{C}}{\mathbb{P}}^{2}.$ We will later see that it is isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}$.
$\bullet$ $\{x_{0}x_{3}-x_{1}x_{2}=0\}\subset{\mathbb{C}}{\mathbb{P}}^{3}.$ This one is isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}\times{\mathbb{C}}{\mathbb{P}}^{1}$.
$\bullet$ $\{x_{0}x_{2}-x_{1}^{2}=x_{0}x_{3}-x_{1}x_{2}=x_{1}x_{3}-x_{2}^{2}=0\}\subset{\mathbb{C}}{\mathbb{P}}^{2}.$ Twisted cubic, isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}$.
$\bullet$ $\{y^{2}-4x^{3}-g_{2}\,x\,z^{2}-g_{3}\,z^{3}=0\}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with coordinates $(x:y:z).$ Elliptic curve.
$\bullet$ $\{x^{3}y+y^{3}z+z^{3}x=0\}\subset{\mathbb{C}}{\mathbb{P}}^{2}.$ Klein quartic; it has $168$ automorphisms.
@@
It remains to argue that $K$ and $C$ are finitely generated modules over ${\mathbb{C}}[x_{0},\ldots,x_{n-1}]$ and apply induction on $n$ (with $n=0$ case just graded finite dimensional vector spaces) to see that $f(t)$ is a rational function with denominator that is a power of (1t)(1-t). The relation to $\dim Z(I)$ is significantly more difficult to prove and the interested reader is referred to [2] or [12].
@@
To find a submodule of $M$ isomorphic to $(A/p)(d)$, let us ask ourselves what this means. Such isomorphism would send $1\in(A/p)(d)$ to some element $m\in M$. The degree of $1$ in $(A/p)(d)$ is, by the definition of the grading shift,5.15.15.1Never trust any signs in any mathematical text ever! equal to (d)(-d), and the ideal $p$ can be recovered as
$p={\rm Ann}(m):=\{a\in A,{\rm~such~that~}a\thinspace m=0\}.$
@@
So now we are simply looking for a nonzero homogeneous element $m\in M$ whose annihilator is prime. As with many things in mathematics, we will try to find such $m$ by imposing on it some maximum or minimum property. As it happens, the right thing to consider is nonzero homogeneous $m$ such that ${\rm Ann}(m)$ is maximum among all such annihilators, in the sense that there are no nonzero homogeneous $m^{\prime}$ with ${\rm Ann}(m^{\prime})\supsetneq{\rm Ann}(m)$. Existence of such $m$ is assured by $A$ being Noetherian. Now suppose that the ideal ${\rm Ann}(m)$ is not prime. It means that there exist $a,b\in A$ with $a\thinspace b\thinspace m=0$, $a\thinspace m\neq 0$, $b\thinspace m\neq 0$. Then ${\rm Ann}(b\thinspace m)$ contains ${\rm Ann}(m)$ and it is strictly larger because it contains $a$ while ${\rm Ann}(m)$ does not. This finishes the proof. ∎
@@
$0\to{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1})(-\deg f_{2})\stackrel{{\scriptstyle f_{2}}}{{\longrightarrow}}{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1})\to{\mathbb{C}}[x_{0},x_{1},x_{2}]/(f_{1},f_{2})\to 0.$
@@
In particular, $S$ contains $x_{0}$, and if we localize ${\mathbb{C}}[x_{0},x_{1},x_{2}]$ at it first we obtain ${\mathbb{C}}[\frac{x_{1}}{x_{0}},\frac{x_{2}}{x_{0}},x_{0}^{\pm 1}]$ (Laurent polynomials in $x_{0}$ over the polynomial ring ${\mathbb{C}}[y_{1},y_{2}]$ for $y_{i}=\frac{x_{i}}{x_{0}}$). When we localize further by the image of $S$ in ${\mathbb{C}}[\frac{x_{1}}{x_{0}},\frac{x_{2}}{x_{0}},x_{0}^{\pm 1}]$ we end up with the $\mathbb{Z}$}-graded ring $R[x_{0}^{\pm 1}]$ where the ring $R$ is defined in (4.2). If we localize the module $M$, we end up with $(R/(\tilde{f}_{1},\tilde{f}_{2}))[x_{0}^{\pm 1}]$ where $\tilde{f}_{i}=\frac{f_{i}}{x_{0}^{\deg f_{i}}}$ is the dehomogenization of $f_{i}$. We can then take degree 0 part of this space to get exactly $R/(\tilde{f}_{1},\tilde{f}_{2})$ that was used to define the multiplicity of the intersection at the point $(1:0:0)$.
So we will take the filtration of $M$ from Proposition 5.1 and observe that the operations of taking localization in $S$ and degree 0 parts preserves the filtration property, with the associated graded objects changing accordingly. If we localize $M={\mathbb{C}}[x_{0},x_{1},x_{2}]/(x_{0},x_{1},x_{2})$ in $S$ we get 0, because $x_{0}$ both annihilates $M_{S}$ and is invertible. Similarly, localization of ${\mathbb{C}}[x_{0},x_{1},x_{2}]/p$ for a prime ideal that corresponds to a point $p$ other than $(1:0:0)$ is zero because there is an element in $S$ which is zero at $p$ but not at $(1:0:0)$ and is therefore both zero and invertible on the localization. Finally, $({\mathbb{C}}[x_{0},x_{1},x_{2}]/(x_{1},x_{2}))_{S}={\mathbb{C}}[x_{0}^{\pm 1}]$ and the degree zero of any grading shift of it is a dimension one vector space over $\mathbb{C}$}. Thus, the number of occurrences of the ideal of $(1:0:0)$ in the filtration is exactly equal to the dimension of $R/(\tilde{f}_{1},\tilde{f}_{2})$ which is the multiplicity of the intersection at this point by (4.3). The same is true for all other points in ${\mathbb{C}}{\mathbb{P}}^{2}$, which finishes the proof of Bezout’s theorem.
@@
$\begin{array}[]{c}\pi^{-1}(U)\enskip\cong\enskip{\mathbb{C}}^{r}\times U\\ \searrow\qquad\swarrow\\ U\end{array}$ (6.4)
@@
Picard group ${\rm Pic}(X)$ is the (abelian) group of line bundles on $X$, up to isomorphism (as line bundles). The product is $\otimes$, the identity is the trivial line bundle ${\mathbb{C}}\times X\to X$, and the inverse is $W\mapsto W^{\vee}$.
@@
The simplest example of a section is the zero section. This means that in every fiber we pick the well-defined element 0 of the corresponding vector space structure. For some line bundles this is the only (holomorphic) section.
@@
We will also introduce some natural bundles on smooth manifolds which will be used later. Namely, for any smooth manifold $X$ we can construct its holomorphic tangent bundle $TX\to X$. Sections of this bundle on $U$ are the same as vector fields on $U$. Its dual $TX^{\vee}\to X$ is called the cotangent bundle. Sections of the cotangent bundle over $U$ are precisely the holomorphic $1$-forms on UU.6.36.36.3When I was learning calculus, the symbol $dx$ really confused me. Was it a function of two variables? Of one variable? I only found peace when it became clear that $dx$ is a section of the cotangent bundle.
@@
Exercise 3. Verify that sections of the line bundle ${\mathbb{C}}\times X\to X$ are in natural bijection to holomorphic functions on $X$.
@@
and note that sections $x_{i}$ of $\mathcal{O}(1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ pull back to holomorphic sections $s_{i}=f^{*}x_{i}$ of $f^{*}\mathcal{O}(1)\to X$. An important observation is that for every point $p\in X$ at least one of the sections $f^{*}x_{i}$ does not equal zero in the fiber of $f^{*}\mathcal{O}(1)\to X$.7.17.17.1We can not make sense of a value of a section, because identification of the fiber with $\mathbb{C}$} is not canonical, but we can ask whether the value is zero, since the fiber has a natural vector space structure. Indeed, the fiber of $f^{*}\mathcal{O}(1)\to X$ at $p$ is the fiber of $\mathcal{O}(1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ at $f(p)$ and we know that not all $x_{i}$ are zero at $f(p)$.
@@
Note that this map is well-defined and holomorphic. The easiest way to see both is to appeal to a local biholomorphism with ${\mathbb{C}}\times U$. While not unique, it is defined up to a holomorphic invertible scaling of the fiber, so all $s_{i}(p)$ are multiplied by the same factor.
@@
As with the Weil divisors, there is a concept of a principal Cartier divisor. It is simply given by $(X,f)$ for a rational function $f$. Clearly, these are identified with principal Weil divisors. In short, we have the following diagram of abelian groups8.68.68.6I have debated whether to write $1$ or 0 on the left, but chose $1$ to honor the multiplicative nature of ${\rm InvRat}(X)$.
$\begin{array}[]{c}1\to{\rm InvRat}(X)\to{\rm Rat}(X)^{*}\to{\rm CartierDiv}(X)\to{\rm CaCl}(X)\to 0\\ \Arrowvert\hskip 60.0pt\Arrowvert\hskip 70.0pt\downarrow\hskip 60.0pt\downarrow\\ 1\to{\rm InvRat}(X)\to{\rm Rat}(X)^{*}\to\quad{\rm WeilDiv}(X)\quad\to\quad{\rm Cl}(X)\quad\to 0\end{array}$ (8.2)
@@
Show that the ideal $I=(x_{1},x_{3})\subset R$ is not principal.8.88.88.8This is an infinitesimal version of the statement that the Weil divisor on the surface $x_{1}x_{2}=x_{3}^{2}$ given by $x_{1}=x_{3}=0]$ can not be locally given as ${\rm div}f$. Hint: Look at the dimension of $I/(x_{1},x_{2},x_{3})I$.
@@
Let $X$ be a smooth complex manifold. Recall from Section 6 that a line bundle $\pi:L\to X$ is a holomorphic map from a smooth complex manifold $L$, with the fibers $\pi^{-1}(x)$ given a structure of a one-dimensional complex vector space. The simplest example is that of a trivial line bundle ${\mathbb{C}}\times X\to X$ and all line bundles are glued from ${\mathbb{C}}\times U\to U$ for some open subsets $U\subseteq X$.
Recall that holomorphic sections of $\pi:L\to X$ are holomorphic maps $s:X\to L$ such that $\pi\circ s={\rm id}_{X}$. If $L={\mathbb{C}}\times X\to X$ is a trivial line bundle, then holomorphic sections of it are in bijection with holomorphic functions $X\to{\mathbb{C}}$. Similarly, we can talk about meromorphic sections of line bundles. Indeed, locally every line bundle is (non-canonically) isomorphic to ${\mathbb{C}}\times U\to U$ and we would be looking at meromorphic functions on $U$ (ratios of two holomorphic functions). If we change to another identification of $\pi^{-1}U$ with ${\mathbb{C}}\times U$, we simply multiply by an invertible holomorphic function, so meromorphicity of sections is well defined.
@@
The identity of the group structure is given by the trivial line bundle ${\mathbb{C}}\times X\to X$ with the section $x\mapsto(1,x)$, and it is easy to check group properties. When checking that the map is a homomorphism, we can work on a cover $X=\bigcup_{\alpha}U_{\alpha}$ for which both $L_{1}$ and $L_{2}$ have trivializations. Then tensor product of sections gives rise to the product of the corresponding functions $f_{\alpha,1}$ and $f_{\alpha,2}$, which corresponds to the product of the Cartier divisors. ∎
@@
Suppose we are given a Cartier divisor $(U_{\alpha},f_{\alpha})_{\alpha\in I}$. Take a copy of ${\mathbb{C}}\times U_{\alpha}$ for each $\alpha$ and glue them together by the following identification. We consider the equivalence relation of the set $\bigsqcup_{\alpha}({\mathbb{C}}\times U_{\alpha})$ with
$({\mathbb{C}}\times U_{\alpha})\ni(t,x)\sim\left(\frac{f_{\beta}(x)}{f_{\alpha}(x)}\,t,x\right)\in({\mathbb{C}}\times U_{\beta})$
@@
For a rational function $f$ on $X$, the principal Cartier divisor ${\rm div}f$ corresponds to the trivial line bundle ${\mathbb{C}}\times X\to X$ with the rational section identified with $f$, i.e.
@@
We can easily verify that the above description fits the definition of Cartier divisor. For example, the ratio $\frac{f_{1}}{f_{2}}=\frac{x_{2}}{x_{1}}$ is an invertible holomorphic function on $U_{1}\cap U_{2}$. The corresponding Weil divisor is built from the zeros and poles of $f_{i}$ and is the line $\{x_{0}=0\}]$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, with coefficient $1$.
@@
This gives us $f_{1}=\left(\frac{x_{0}}{x_{1}}\right)^{-3}$ (signs, or more generally invertible functions do not matter) and similarly $f_{2}=\left(\frac{x_{0}}{x_{2}}\right)^{-3}$. Then the corresponding Weil divisor is $3[\{x_{0}=0\}]$, and therefore the canonical line bundle over ${\mathbb{C}}{\mathbb{P}}^{2}$ is isomorphic to $\mathcal{O}(-3)$.
@@
[…]
@@
$\alpha_{2}$$\alpha_{1}\infty$$\alpha_{3}$$\alpha_{2}$$\alpha_{1}\infty$$\alpha_{3}$
@@
for the newly defined addition operation ++. We claim that it gives $E$ a group structure, which is of course abelian by construction, since lines $AB$ and $BA$ are the same.
@@
$z_{0}$$z_{0}+\gamma_{1}$$z_{0}+\gamma_{2}$$z_{0}+\gamma_{1}+\gamma_{2}$$D$
The sum of residues in $D$ is the counterclockwise integral $\int_{\partial D}f(z)\,dz$. By periodicity of $f$, the contributions of the opposite sides of $D$ cancel. ∎
@@
$\frac{1}{2\pi{\rm i}}\int_{z=z_{0}}^{z=z_{0}+\gamma_{1}}z\frac{f^{\prime}(z)}{f(z)}\,dz-\frac{1}{2\pi{\rm i}}\int_{z=z_{0}+\gamma_{2}}^{z=z_{0}+\gamma_{1}+\gamma_{2}}z\frac{f^{\prime}(z)}{f(z)}\,dz$
$=\frac{1}{2\pi{\rm i}}\int_{z=z_{0}}^{z=z_{0}+\gamma_{1}}\left(z\frac{f^{\prime}(z)}{f(z)}-(z+\gamma_{2})\frac{f^{\prime}(z+\gamma_{2})}{f(z+\gamma_{2})}\right)\,dz=-\frac{1}{2\pi{\rm i}}\int_{z=z_{0}}^{z=z_{0}+\gamma_{1}}\gamma_{2}\frac{f^{\prime}(z)}{f(z)}\,dz$
@@
We know that the residue at $z=0$ must be 0, so our first attempt at finding such $f$ is
@@
Great, this looks periodic. But, we have a slight problem – the series does not have absolute convergence. Indeed, the number of lattice points $l$ of size approximately $k$ is on the order of ckc\,k for some constant $c$, so we have roughly ckc\,k terms of size $\frac{1}{k^{2}}$, and the harmonic series diverges (although rather slowly).
@@
In the last equality we used that the change of variables $l^{\prime}\to-m-l^{\prime}$ sends $S=\sum_{l^{\prime}\neq-m,0}\Big(\frac{1}{(l^{\prime})^{2}}-\frac{1}{(l^{\prime}+m)^{2}}\Big)$ to (S)(-S), which implies $S=0$.
@@
For every $a\not\in\frac{1}{2}L$ the even elliptic function $\mathcal{P}(z)-\mathcal{P}(a)$ has zeros at $a\,{\rm mod}\,L$ and $a\,{\rm mod}\,L$ with multiplicity one, and no other zeros. For $a\in\frac{1}{2}L\setminus L$ the function $\mathcal{P}(z)-\mathcal{P}(a)$ has zeros of order two at $a\,{\rm mod}\,L$ and no other zeros.
@@
The function $\mathcal{P}(z)-\mathcal{P}(a)$ has a pole of order two at $L$ and no other poles. Therefore, by Propositions 11.3 and 11.4 it has two zeros (with multiplicity, up to $L$) which add up to $0\,{\rm mod}\,L$. One of these zeros is $a$, so the other is (a)(-a). The situation with $2a\in L$ is special, because then we have $a=-a\,{\rm mod}\,L$ and a double zero. ∎